# The Audio-Visual Conversational Graph: From an Egocentric-Exocentric Perspective

### v1

@@ Abstract
In recent years, the thriving development of research related to egocentric videos has provided a unique perspective for the study of conversational interactions, where both the visual and audio signals play a crucial role. While most prior work focus on learning about behaviors that directly involve the camera wearer, we introduce the Ego-Exocentric Conversational Graph Prediction problem, marking the first attempt to infer exocentric conversational interactions from egocentric videos. We propose a unified multi-modal, multi-task framework — Audio-Visual Conversational Attention (AV-Conv), for the joint prediction of conversation behaviors—speaking and listening—for both the camera wearer as well as all other social partners present in the egocentric video. Specifically, we customize the self-attention mechanism to model the representations across-time, across-subjects, and across-modalities. To validate our method, we conduct experiments on a challenging egocentric video dataset that includes first-person perspective, multi-speaker, and multi-conversation scenarios. Our results demonstrate the superior performance of our method compared to a series of baselines. We also present detailed ablation studies to assess the contribution of each component in our model. footnotetext: * This work was done during an internship at Reality Labs Research.
@@ 1 Introduction
Figure 1: We propose (d) the Ego-Exocentric Conversational Graph Prediction problem that jointly learns (a) the egocentric behaviors—whether the camera wearer is speaking or listening to others), and (b) the exocentric behaviors—whether the other social partners in the scene are speaking or listening to one another, given only the egocentric video input (c).
@@
Motivated by the above, we introduce the concept of Audio-Visual Conversational Graph, which describes the conversational behaviors—speaking and listening—for both the camera wearer and all social partners involved in the conversation. As shown in Fig. 1, we present a challenging Ego-Exocentric Conversational Graph Prediction problem. When provided with an egocentric video containing multiple people actively engaged in a conversation, our goal is to create a complete dynamic directed graph that can instantly reflect the conversational behaviors and relationships among all participants. Unlike prior work that focuses on only a single task [18, 34], we simultaneously address multiple closely related tasks in a multi-task setting, and we are the first to explicitly predict the dense conversational behaviors from an exocentric point of view.
@@
  • We propose a unified multi-modal, multi-task framework that jointly predicts the interaction states of all social entities captured in the egocentric videos.
@@ 3.2 Dataset and Annotations
Instead, we make use of the recently introduced Egocentric Concurrent Conversations Dataset [34], which contains a total of 50 participants, evenly distributed across 10$\sim$30-minute data collection sessions, with each session comprising groups of five individuals. Each individual wears a headset with an Intel SLAM camera and an array of six microphones during the sessions, resulting in $\sim$20 hours of egocentric videos in total. This means in each session with five people in total, each of them serves as the camera wearer in their egocentric video, and the maximum number of the other social partners captured in the egocentric video frames is four.
@@ 4 The Audio-Visual Conversational Attention Model
In the following, we first explain how we extract audio-visual representations for analyzing conversational behaviors (Sec. 4.1). Then, we present our Audio-Visual Conversational Attention (AV-Conv) model that disentangles the correlations of the representations from different subjects with an attention mechanism (Sec. 4.2). Finally, we introduce the classifiers used to predict the edge attributes (Sec. 4.3).
@@ 4.1 Audio-Visual Feature Extraction
Below we introduce the representations we use for the location of all speakers, their visual appearance, and the multi-channel audio.
Speaker Location. To efficiently capture the location information of each individual, we use the head bounding boxes annotations as mentioned in Sec. 3.2 to crop all visible heads in the video frames. Similar to [34], we use binary masks at the locations of the cropped heads, denoted as $\mathbf{S}\in\mathbb{R}^{N\times 1\times T\times h\times w}$, to preserve the spatial positions for all social partners.
@@ 4.2 Conversational Attention
Cross-Subject Attention. Our task demands a simultaneous prediction of diverse conversational behaviors for all subjects, thus making it essential to allow representations of each subject to attend to each other for understanding the interpersonal dynamics. $\mathcal{T}_{N}$ takes in token $\mathbf{Z}^{av}_{t}$ and applies self-attention to each person’s spatiotemporal patch feature over the subject dimension. In this way, we enable the interactions of the audio-visual spatiotemporal features of each individual, thereby increasing the distinction in predictions across different subjects. It outputs token $\mathbf{Z}^{av}_{tn}$.
Global-Local Attention. As depicted in Section 4.1, the head feature is derived from each individual captured in the frame, while the audio feature encompasses both global information about the entire scene and local positional information of the faces. We also apply cross-modality Global-Local self-attention to compute the correlations between the global audio-positional tokens and the local head appearance tokens, enhancing the awareness of the entire scene for each feature and the semantic alignment. $\mathcal{T}_{S}$ takes in token $\mathbf{Z}^{av}_{tn}$ and compares all patches with each other.
A residual connection is used to aggregate output from each self-attention layer, and the final feature passes through another fully-connected layer. Overall, with a given audio-visual input, we denote the augmented feature $\mathbf{Z}^{av}_{tns}$ as:
@@ 4.3 Conversation Edge Attributes Classification
Method Egocentric Graph Exocentric Graph
$e^{S}_{c\rightarrow p_{i}}$ $e^{S}_{p_{i}\rightarrow c}$ $e^{L}_{c\rightarrow p_{i}}$ $e^{L}_{p_{i}\rightarrow c}$ $e^{S}_{p_{i}\rightarrow p_{j}}$ $e^{S}_{p_{j}\rightarrow p_{i}}$ $e^{L}_{p_{i}\rightarrow p_{j}}$ $e^{L}_{p_{j}\rightarrow p_{i}}$
SAAL (Acc) 86.23 67.10 86.48 86.99 / / / /
ASL+Layout (Acc) 13.71 55.53 83.48 77.35 65.96 43.63 64.49 79.04
AV-Conv (Acc) 90.02 75.94 87.80 90.63 86.15 75.89 75.91 85.75
SAAL (mAP) 68.43 44.97 44.64 39.55 / / / /
ASL+Layout (mAP) 86.28 47.45 21.83 47.91 45.91 46.68 18.98 16.15
AV-Conv (mAP) 82.08 68.94 60.70 65.48 72.73 63.36 32.35 29.29
@@ 5.2 Baselines
Since we are the first to address the challenging problem of conversational graph, there are no pre-existing baselines directly comparable. We compare with two baselines adapted from prior work [34, 18], and a series of ablated versions of our method.
@@ 5.4 Visualization
In Fig. 5.3, we visualize some examples from the Egocentric Concurrent Conversations Dataset and show the prediction results of the conversational graph from our model. We can see that it is a very challenging problem, as the subjects in the video frames are all visually similar and exhibit complex conversational interactions—there can be multiple people speaking and listening at the same time. Despite the challenges and diverse people layout, our model makes accurate predictions of the complete conversation behaviors for both the camera wearer as well as all other social partners present in the scene from just the egocentric videos, demonstrating the effectiveness of our AV-Conv model.
@@ 6 Conclusion
We presented the Audio-Visual Conversational Graph Prediction problem to infer exocentric conversational interactions from egocentric videos. With our unified multi-modal, multi-task framework AV-Conv, we have demonstrated the superiority of our proposed method on a challenging multi-speaker, multi-conversation egocentric dataset. Our work marks a new research direction to learn exocentric social interactions from egocentric videos. As future work, we plan to extend our framework to jointly solve other social behavior tasks such as gaze prediction, and study more complex social relationships such as conversation groups.

### v2

@@ Abstract
In recent years, the thriving development of research related to egocentric videos has provided a unique perspective for the study of conversational interactions, where both visual and audio signals play a crucial role. While most prior work focus on learning about behaviors that directly involve the camera wearer, we introduce the Ego-Exocentric Conversational Graph Prediction problem, marking the first attempt to infer exocentric conversational interactions from egocentric videos. We propose a unified multi-modal framework—Audio-Visual Conversational Attention (AV-Conv), for the joint prediction of conversation behaviors—speaking and listening—for both the camera wearer as well as all other social partners present in the egocentric video. Specifically, we adopt the self-attention mechanism to model the representations across-time, across-subjects, and across-modalities. To validate our method, we conduct experiments on a challenging egocentric video dataset that includes multi-speaker and multi-conversation scenarios. Our results demonstrate the superior performance of our method compared to a series of baselines. We also present detailed ablation studies to assess the contribution of each component in our model. Check our Project Page. footnotetext: * This work was done during an internship at Meta Reality Labs.
@@ 1 Introduction
Figure 1: We propose (d) the Ego-Exocentric Conversational Graph Prediction problem that jointly learns (a) the egocentric behaviors—whether the camera wearer is speaking or listening to others, and (b) the exocentric behaviors—whether the other social partners in the scene are speaking or listening to one another, given only the egocentric video input (c).
@@
Motivated by the above, we introduce the concept of Audio-Visual Conversational Graph, which describes the conversational behaviors—speaking and listening—for both the camera wearer and all social partners involved in the conversation. As shown in Fig. 1, we present a challenging Ego-Exocentric Conversational Graph Prediction problem. When provided with an egocentric video containing multiple people actively engaged in a conversation, our goal is to create a complete dynamic directed graph that can instantly reflect the conversational behaviors and relationships among all participants. Unlike prior work that focuses on only a single task [19, 36, 18], we simultaneously address multiple closely related tasks in a unified framework, and we are the first to explicitly predict the dense conversational behaviors from an exocentric point of view.
@@
  • We propose a unified multi-modal framework that jointly predicts the interaction states of all social entities captured in the egocentric videos.
@@ 3.2 Dataset and Annotations
Instead, we make use of the recently introduced Egocentric Concurrent Conversations Dataset [36], which contains a total of 50 participants, evenly distributed across 10$\sim$30-minute data collection sessions, with each session comprising groups of five individuals. Each individual wears a headset with an Intel SLAM camera and an array of six microphones during the sessions, resulting in $\sim$20 hours of egocentric videos in total. This means in each session with five people in total, each of them serves as the camera wearer in their egocentric video, and the maximum number of the other social partners captured in the egocentric video frames is four. This dataset setting aligns with the prior psychological study [22] which reveals a size constraint of about four people for meaningful group conversations.
@@ 4 The Audio-Visual Conversational Attention Model
In the following, we first explain how we extract audio-visual representations (Sec. 4.1). Then, we present our Audio-Visual Conversational Attention (AV-Conv) model that disentangles the correlations of the representations from different subjects with an attention mechanism (Sec. 4.2). Finally, we introduce the classifiers used to predict the edge attributes (Sec. 4.3).
@@ 4.1 Audio-Visual Feature Extraction
Below we introduce representations of the location of all speakers, their visual appearance, and multi-channel audio.
Speaker Location. To efficiently capture the information of each individual, we use the head bounding boxes annotations mentioned in Sec. 3.2 to crop all visible heads in the video frames. We also follow  [36] to use binary masks at the locations of the cropped heads, denoted as $\mathbf{S}\in\mathbb{R}^{N\times 1\times T\times h\times w}$, to preserve the spatial positions for all social partners.
@@ 4.2 Conversational Attention
Cross-Subject Attention. Our task demands a simultaneous prediction of diverse conversational behaviors for all subjects, thus making it essential to allow representations of each subject to attend to each other for understanding the interpersonal dynamics. $\mathcal{T}_{N}$ processes token $\mathbf{Z}^{av}_{t}$, applying self-attention to tokens at corresponding spatiotemporal locations for each person. This facilitates interactions among audio-visual features of individuals, enhancing prediction distinctions across subjects. It yields token $\mathbf{Z}^{av}_{tn}$.
Global-Local Attention. As depicted in Section 4.1, the head feature is derived from each individual in the frame, and the audio feature contains both scene-wide details and facial positional information. We also apply cross-modality Global-Local self-attention to correlate global audio-positional tokens with local head appearance tokens, enhancing the scene awareness and semantic alignment for each feature. $\mathcal{T}_{S}$ takes in token $\mathbf{Z}^{av}_{tn}$ and compares all patches with each other.
A residual connection aggregates output from each self-attention layer, and the final feature passes through another fully connected layer. Overall, with a given audio-visual input, we denote the augmented feature $\mathbf{Z}^{av}_{tns}$ as:
@@ 4.3 Conversation Edge Attributes Classification
Method Egocentric Graph Exocentric Graph Ego-Exo Avg
$e^{S}_{c\rightarrow p_{i}}$ $e^{S}_{p_{i}\rightarrow c}$ $e^{L}_{c\rightarrow p_{i}}$ $e^{L}_{p_{i}\rightarrow c}$ $e^{S}_{p_{i}\rightarrow p_{j}}$ $e^{S}_{p_{j}\rightarrow p_{i}}$ $e^{L}_{p_{i}\rightarrow p_{j}}$ $e^{L}_{p_{j}\rightarrow p_{i}}$ $G_{conv}$
SAAL (Acc) 86.23 67.10 86.48 86.99 / / / / /
ASL+Layout (Acc) 13.71 55.53 83.48 77.35 65.96 43.63 64.49 79.04 60.40
AV-Conv (Acc) 90.02 75.94 87.80 90.63 86.15 75.89 75.91 85.75 83.51
SAAL (mAP) 68.43 44.97 44.64 39.55 / / / / /
ASL+Layout (mAP) 86.28 47.45 21.83 47.91 45.91 46.68 18.98 16.15 41.40
AV-Conv (mAP) 82.08 68.94 60.70 65.48 72.73 63.36 32.35 29.29 59.37
@@ 5.2 Baselines
Since we are the first to address the challenging problem of conversational graph,there are no pre-existing baselines directly comparable. We adapt two baselines from prior work [36, 19], and various ablated versions of our method.
@@ 5.4 Visualization
In Fig. 5, we visualize some examples from the Egocentric Concurrent Conversations Dataset and show the prediction results of the conversational graph from our model. We can see that it is a very challenging problem, as the subjects in the video frames are all visually similar and exhibit complex conversational interactions—there can be multiple people speaking and listening at the same time. Despite the challenges and diverse people layout, our model makes accurate predictions of the complete conversation behaviors for both the camera wearer as well as all other social partners present in the scene from just the egocentric videos, demonstrating the effectiveness of our AV-Conv model.
@@ 6 Conclusion
We presented the Audio-Visual Conversational Graph Prediction task to infer exocentric conversational interactions from egocentric videos. Our AV-Conv model shows superiority on a challenging multi-speaker, multi-conversation egocentric dataset, paving the way for learning exocentric social interactions from egocentric videos. Future plans involve extending our framework to tackle additional social behavior tasks like gaze prediction and exploring more intricate social relationships within conversation groups.